# arXiv:2602.24001v1[math.DS] 27 Feb 2026

### v1

@@ Abstract.
The so-called Filament Based Lamellipodium Model is a complex modeling framework for a very heterogeneous chemo-mechanical system of cell biology. It contains a model for Coulomb repulsion between filaments, whose effect on the stability of the system has been unclear. In this work, a strongly simplified version of the model is considered, showing a destabilizing effect of the repulsion, leading to a pitchfork bifurcation from a trivial steady state to traveling wave solutions. The simplified model is derived, its linearization around the trivial steady state is analyzed, and a formal bifurcation analysis is carried out. It is shown that the pitchfork bifurcation maybe super- or sub-critical. Time dependent numerical simulations illustrate these results and provide additional, more global information.
@@ 1. Introduction
In Section 3, the linearization around the trivial steady state is analyzed. The essential part of the linearized problem is a cross-diffusion system, which is shown to be well posed. The null space of the linearized operator is generically one-dimensional due to the rotational symmetry. However, a steady state bifurcation occurs for a critical value of a dimensionless parameter, measuring the combined strength of the inward forces and the tension relative to the strength of the pressure. For large enough values of this parameter, spectral stability of the set of trivial steady states is shown.
These results raise the expectation that the trivial steady states can be destabilized by the effect of the pressure. This is verified in Section 4 by a formal bifurcation analysis. Since the problem has an additional flip symmetry, a pitchfork bifurcation occurs, which may be super- or sub-critical, depending on the value of an additional dimensionless parameter. As a consequence of the rotational symmetry, the computations are rather involved and therefore aided by using Mathematica [11]. The resulting normal form is a set of two ODEs. One of them has the standard normal form of the pitchfork bifurcation, whereas the second unknown passively follows the dynamics of the first. The bifurcating solutions are only stationary concerning the shape of the filament ensemble which is deformed compared to the trivial steady state. Since in the deformed states the outer filaments are not parallel, the ensemble moves with constant velocity, pushed by the inward forces.
In the last section, the bifurcation results are illustrated by numerical simulations of the time dependent problem. A finite volume discretization is used for the spatial variable and a standard ODE solver for the time discretization. Typical solutions for parameter values before and after the bifurcation are computed. In the case of the subcritical bifurcation, the simulation shows convergence to a deformed state far from the trivial equilibrium. This suggests that the bifurcating branch, which is initially unstable, undergoes a secondary fold bifurcation.
@@ 4. Normal form reduction close to the bifurcation point
Table 1. Behavior of steady states depending on $\beta$ and $\gamma$
@@ 6. Discussion
Our formal bifurcation analysis and numerical simulations reveal that this instability is the normal form of a pitchfork bifurcation. The nature of this bifurcation, whether super- or sub-critical, depends on the tension parameter $\gamma$. The sub-critical case is of particular interest mathematically as it implies the existence of a hysteresis loop where the system may jump to a large-deformation branch even before the linear stability threshold is crossed. This suggests that in certain parameter regimes, the filament network is highly sensitive to perturbations, potentially allowing the system to switch between stable symmetric states and deformed states.

### v2

# Pitchfork bifurcation and traveling waves for a planar ensemble of rigid filaments with repulsive interaction

@@ Abstract.
The so-called Filament Based Lamellipodium Model is a complex modeling framework for a very heterogeneous chemo-mechanical system of cell biology. It contains a model for Coulomb repulsion between filaments, whose effect on the stability of the system has been unclear. In this work, a strongly simplified version of the model is considered, showing a destabilizing effect of the repulsion. This instability results in a pitchfork bifurcation with an additional rotational symmetry, leading to a two-dimensional bifurcating manifold of traveling wave solutions. The simplified model is derived, its linearization around the trivial steady state is analyzed, and a formal bifurcation analysis is carried out. It is shown that the pitchfork bifurcation maybe super- or sub-critical. Time dependent numerical simulations illustrate these results and provide additional, more global information on the emergence of periodic and chaotic dynamics by secondary bifurcations.
@@
Corresponding author
Abbreviated title: Bifurcation and waves in repulsive filaments
@@ 1. Introduction
In Section 3, the linearization around the trivial steady state is analyzed. The essential part of the linearized problem is a cross-diffusion system, which is shown to be well posed. The null space of the linearized operator is generically one-dimensional due to the rotational symmetry. However, a steady state bifurcation occurs for a critical value of a dimensionless parameter, measuring the combined strength of the inward forces and the tension relative to the strength of the pressure. For large enough values of this parameter, the spectral stability of the set of trivial steady states is shown.
These results raise the expectation that the trivial steady states can be destabilized by the effect of the pressure. This is verified in Section 4 by a formal bifurcation analysis. Since the problem has an additional flip symmetry, a pitchfork bifurcation occurs, which may be super- or sub-critical, depending on the value of an additional dimensionless parameter. Crucially, due to the system’s inherent rotational symmetry, we do not just have a standard pitchfork bifurcation; rather, the bifurcating manifold is two-dimensional. As a consequence of this symmetry, the computations are rather involved and, therefore, aided by using Mathematica [11]. The resulting normal form is a set of two ODEs. One of them governs the amplitude of the bifurcation, whereas the second unknown passively follows the dynamics of the first to account for the rotational invariance. The bifurcating solutions are stationary only in relation to the shape of the filament ensemble, which is deformed compared to the trivial steady state. Since in the deformed states the outer filaments are not parallel, the ensemble moves with constant velocity, pushed by the inward forces.
In the last section, the bifurcation results are illustrated by numerical simulations of the time dependent problem. A finite volume discretization is used for the spatial variable and a standard ODE solver for the time discretization. Typical solutions for parameter values before and after the bifurcation are computed. In the case of the subcritical bifurcation, the simulation shows convergence to a deformed state far from the trivial equilibrium. This suggests that the bifurcating branch, which is initially unstable, undergoes a secondary fold bifurcation. Furthermore, beyond these traveling wave solutions, our numerical simulations reveal a much richer dynamical landscape, including the emergence of both periodic and chaotic dynamics within the network.
@@ 4. Normal form reduction close to the bifurcation point
Table 1. Behavior of steady states depending on the dimensionless parameters $\beta$ and $\gamma$
@@ 6. Discussion
Our formal bifurcation analysis and numerical simulations reveal that this instability is the normal form of a pitchfork bifurcation. However, owing to the rotational symmetry of the system, this is not a standard one-dimensional pitchfork, instead, the bifurcating manifold is two-dimensional. The nature of this bifurcation, whether super- or sub-critical, depends on the tension parameter $\gamma$. The sub-critical case is of particular interest mathematically as it implies the existence of a hysteresis loop where the system may jump to a large-deformation branch even before the linear stability threshold is crossed. Beyond transitioning to steadily moving deformed states, our numerical investigations demonstrate that the system also exhibits more complex behaviors, including periodic oscillations and chaotic dynamics. This suggests that in certain parameter regimes, the filament network is highly sensitive to perturbations, potentially allowing the system to switch between stable symmetric states, persistent traveling waves, and even chaotic regimes.